# A Comprehensive Survey on Instruction Following

### v7

@@ 5.1 Semantic Parser
At the early stage of machine learning, to help the systems understand natural language instructions, a great number of works employed semantic parsing to convert the instruction into the formal language (logical formula), which can be more easily executed by the systems (Goldwasser and Roth, 2014, inter alia). As exemplified in Figure 3 (a), a game instruction “Move any top card to an empty free cell” can be processed into an executable formula: “card(x) $\wedge$ freecell(y)”.
@@ 5.2 Flatten-and-Concatenation
In contrast to the semantic parser approach, which considers the instructions’ structure and the target problems, methods based on the neural networks take more brutal treatment: as illustrated in Figure 3 (b)—instructions, regardless of their length, structure, task types, etc., are flattened as a long token sequence and concatenated with the input X as a new input sequence for the models, which has been widely adopted by the prior research (Wang et al., 2022d; Wei et al., 2023, inter alia). However, this naive strategy constantly results in unsatisfactory performances when using vanilla models Weller et al. (2020), leading to its reliance on large-scale instruction fine-tuning, well-known as “instruction following”.
@@ 5.3 HyperNetwork
Unlike the conventional modeling strategy that encodes the input sequence into the dense representation (i.e., language-to-representation), hypernetwork follows a language-to-parameter paradigm: as shown in Figure 3 (c), this scheme converts textual instruction into a block of model parameters that can be further plugged into the underlying models Ha et al. (2016); Houlsby et al. (2019); Jin et al. (2020). As a result, hypernetwork-based instruction modeling can better leverage the structured input sequence by encoding the instruction and task input separately (i.e., instruction-to-parameter, input-to-representation), achieving stronger generalization compared with the flatten-and-concatenation approach Ye and Ren (2021); Deb et al. (2022). It can also significantly improve inference efficiency, as concluded by recent works Ivison et al. (2022).
@@ 5.4 Reinforcement Learning from Human Feedback
As shown in Figure 3 (d), given the initial LM and its copy (the target tuned LM), for each input sequence (e.g., instruction and task input), these two different LM will generate two outputs. After getting the generation texts of the initial and tuned LM, we can further calculate the textual difference penalty between them:
$y$ $=f_{\theta}(I,x)$
$y^{*}$ $=f_{\theta^{*}}(I,x)$
@@
Here, the $f_{\theta}$ and $f_{\theta^{*}}$ represent the parameters of initial and tuned LM, respectively. $I$ and $x$ denote the instruction and task input. While $y$ and $y^{*}$ are the outputs of the initial and tuned LM. $r_{\textrm{KL}}$ is the final reward (loss) of prediction shifting, and the KL means the calculation of Kullback–Leibler (KL) divergence. KL divergence is a widely adopted strategy for measuring the textual difference, which can be used as a part of the loss to penalize the tuned LM on shifting the output substantially away from the initial LM’s generation. This prediction shift penalty can prevent the tuned LM from fooling the reward function to get a high reward but loses the coherence of the generation text.
@@
The $r_{\theta}$ here is the parameters of the reward model, which is usually a regression model.
@@
Where the $\sigma$ is the activation function that scales the reward into $(0,1]$, $y^{+}$ is the output preferred by human, $y^{-}$ vice versa. By training on these pairwise preference comparison data, the reward model can directly learn to capture the human preference and make alignment reward estimation for the RLHF.
@@ 6.1 Dataset
The essence of instruction following is to tame the models by following various task instructions and responding with the corresponding desired outputs. Therefore, the instruction-tuning dataset (high-quality instruction-output pairs) is the critical part Wang et al. (2023b); Zhou et al. (2023a). The current instruction-tuning datasets can be divided into two categories according to different annotation categories: 1) human-annotated datasets (§ 6.1.1); 2) LLM-synthetic datasets (§ 6.1.2).
@@ 6.1.1 Human-annotated datasets
The conventional way to create instruction-tuning datasets is by employing human annotators, especially for those early-stage datasets, as shown in Table 6. For example, Public Pool of Prompts (P3) (Sanh et al., 2022) and Flan (Wei et al., 2022a) collected multi-task instruction-tuning datasets, where they employed human expertise to design various prompts for each task. Mishra et al. (2022b) proposed Natural Instructions, in which they collected more than 60 NLP tasks with the corresponding human-written instructions; Wang et al. (2022d) further extended this collection into a 1.6k cross-lingual tasks scale contributed by 88 NLP experts, namely Super-Natural Instructions.
@@ 6.1.2 LLM-synthetic datasets
Since LLMs have shown their superior annotation quality on various NLP tasks He et al. (2023); Pan et al. (2023), tons of recent works tried to employ LLMs (e.g., ChatGPT and GPT-4) instead of humans on instruction-tuning dataset curation. For instance, Self-Insturct Wang et al. (2022c) and Unnatural Instructions Honovich et al. (2022a) utilized human-annotated instructions as demonstrations to guide LLMs in devising novel tasks and increasing task diversity. WizardLM Xu et al. (2023a) employed an instruction evolution paradigm to increase instruction complexity. Dynosaur Yin et al. (2023) repurposed existing input-output pairs in NLP datasets to stimulate new instructions and reduce annotation costs. MUFFIN Lou et al. (2023) prompted the LLMs to gather different task instructions for the same input and obtained an impressive generalization capacity of the tuned smaller models. Besides single-turn instruction-output datasets, some works also collected multi-turn dialogue data from ShareGPT [^3] https://sharegpt.com/, where the instructions are created by humans (users of OpenAI API), and the responses are from LLMs Geng et al. (2023); Chiang et al. (2023).
Though these LLM-synthetic datasets contained considerable noise (e.g., incoherent instructions and hallucination outputs), the diverse task distribution and model-preferred output patterns still benefit the smaller models on instruction-following, achieving comparable even better generalization performance compared with human-annotated datasets Wang et al. (2022c, 2023b).
In a word, the choice between human-annotated and LLM-synthetic datasets can also be regarded as a trade-off between data quality and diversity. Previous works have concluded that both factors affect the performance of the resulting models Chung et al. (2022); Longpre et al. (2023) — mixing human and machine data can lead to better results Wang et al. (2022c); Yin et al. (2023), while there is no concrete conclusion about which factor outweighs the other, which highly depends on the downstream tasks and application situations.
@@ 6.2.1 Different evaluation schemes
When testing the model’s instruction-following performance on an evaluation dataset, if this dataset has “ground-truth” outputs, then a conventional criterion is to use those automatic evaluation metrics, such as Exact-Match Rajpurkar et al. (2016) and ROUGE Lin (2004), that have been widely used for evaluating the generation models Mishra et al. (2022b); Wang et al. (2022d); Wei et al. (2022a); Sanh et al. (2022); Lou and Yin (2023); Yin et al. (2022). However, this naive evaluation strategy suffers from several drawbacks: 1) it has been widely committed that the automatic generation metrics are not perfect and have significant biases (e.g., BLUE score has text length bias); 2) all of these metrics are used for showing how well the model’s prediction aligns with pre-annotated answers, however, most real-world user tasks are highly open-ended, and there are probably no official ground-truth labels to calculate the metrics; 3) the essence of instruction following is to follow user’s instructions and provide desired responses that can well address user’s requirements, while automatic metrics more focus on some superficial textual patterns and lacking the reflection on how well the response satisfies the instructions.
@@
A more reliable evaluation method is to employ humans to decide whether a model’s response satisfies the instruction or not. For example, given a task instruction and a corresponding model output, the human evaluator should read the instruction and decide whether this model output is acceptable or not (reporting an acceptance ratio for the target model) Wang et al. (2022c, 2023b); Lou et al. (2023); or ask humans to compare two models’ outputs and decide which one better satisfies the instruction (pairwise comparison between two models) Chiang et al. (2023); Geng et al. (2023); Taori et al. (2023). Since instructions are mostly complicated and contain a lot of explicit or implicit constraints, human evaluation is more flexible and accurate than automatic metrics in reflecting the instruction-following capacities of different models.
However, human evaluation is also much more expensive and slower than automatic evaluation. Thus, most of the works only conduct human evaluation on a small subset of the whole evaluation benchmark. Meanwhile, it is mostly based on human evaluators’ personal preferences and can result in high variance between different evaluators.
@@
To address the aforementioned issues of human evaluation, recent works also tried to use LLMs (e.g., GPT-4) rather than humans to evaluate the models’ instruction following capacity, such as VicunaEval Chiang et al. (2023) and AlpacaEval Taori et al. (2023). Nevertheless, although LLMs are cheaper and faster, they were found to have serious preference bias on some superficial textual patterns or hallucinations, e.g., GPT-4 prefers the longer texts and those responses with diverse tokens Wang et al. (2023b). Meanwhile, only a final preference score is usually insufficient for a comprehensive evaluation.
In order to improve reliability, instead of letting LLMs simply provide a preference decision, other works tend to ask LLMs to generate comprehensive analyses besides the final decision, such as generating the error types, locations and explanations before concluding with the final scores Fernandes et al. (2023); Xu et al. (2023d). Some other works also predefined several explainable criterion questions for the various evaluation tasks (e.g., for an instruction “Please generate at least 25 sentences”, define a criterion “is the model’s generation at least 25 sentences?”), that can be further verified by humans or LLMs easily (i.e., doing binary classification on these predefined criteria) Liu et al. (2023e); Zhou et al. (2023b). Saha et al. (2023) also asked LLMs to first generate the criteria questions automatically according to the instructions and then evaluate the model’s response.
@@ 6.2.2 Two branches of evaluation
Most evaluation datasets in this branch are based on conventional multi-task learning, where the evaluation tasks are mostly traditional NLP tasks, such as natural language inference Wei et al. (2022a); Sanh et al. (2022). This branch aims to test LLMs’ instruction-following and problem-solving capacity, and the main criterion here is whether the models can correctly solve the given textual task. Therefore, most of the evaluation settings in this branch adopt conventional automatic metrics to reflect the task ground-truth label alignment. Representative benchmarks are MMLU Hendrycks et al. (2020), BBH Suzgun et al. (2022), SuperNI-Test Wang et al. (2022d), T0-Eval Sanh et al. (2022), InstructEval Chia et al. (2023), etc.
@@ 7 Factors that Influence Instruction Following Performance
Recipes for Instruction Following
Model-related Factors
(§ 7.1)
@@
Instruction-related Factors
(§ 7.2)
@@
Demonstration-related Factors
(§ 7.3)
@@
Model-Instruction Alignment
(§ 7.4)
@@
Data-wise Factors
(§ 7.5)
@@ 7.1.1 Update model or not
As shown in Figure 1 (b), to drive LLMs to understand and follow task instructions more smoothly, a widely-adopted practice is fine-tuning LLMs on multi-task datasets, where each task input is equipped with a task instruction. This procedure is also well-known as “instruction following”. A lot of works demonstrated that instruction-tuned LLMs could better follow the instructions of unseen tasks compared with frozen LLMs Wei et al. (2022a); Sanh et al. (2022).
Besides the performance gains on unseen tasks, instruction following has many other benefits, such as learning faster on the downstream tasks Longpre et al. (2023); Gupta et al. (2023), being more robust to the tiny instruction perturbations (e.g., paraphrasing) Weller et al. (2020); Sanh et al. (2022); Gu et al. (2022), becoming more user-friendly Chung et al. (2022), and being better at following soft instructions Wei et al. (2022a), etc.
@@ 7.1.2 Model scale
Recent works demonstrated that the model scale significantly impacts the generalization performance of instruction following (Chung et al., 2022; Longpre et al., 2023; Wang et al., 2023b, inter alia). As shown in Figure 4, the generalization performances of each model consistently increase when scaling up the model size. More interestingly, when the model scale is large enough, even vanilla LLMs can significantly outperform smaller LLMs tuned on extensive tasks (see Flan-PaLM; vanilla 540B $>$ 8B + 1836 tasks), which probably implies that the benefits of scaling up the model size can outweigh dataset scaling.
However, the super-large model scale is usually unaffordable for most research groups, and it also leads to huge carbon emissions, making it unrealistic in most real-world scenarios Strubell et al. (2019); Schick and Schütze (2021c). Accordingly, recent works began to investigate a more efficient way to address the model scale problem, e.g., the parameter-efficient fine-tuning Hu et al. (2021); Liu et al. (2022a); Lialin et al. (2023); Jang et al. (2023).
@@ 7.2.1 Instruction engineering
One straightforward solution is to manually rewrite the instruction, i.e., human instruction engineering. When humans perform instruction engineering, the criteria of rewriting are mostly human intuition. For example, Mishra et al. (2022a) conducted error cases analysis on GPT’s instruction-following outputs. Accordingly, they designed several empirical rules on instruction writing and “reframed” the instructions. All of these proposed rules are based on human intuition, e.g., itemizing instructions and task decomposition. In order to avoid the preference bias introduced by a small group of humans, Bach et al. (2022) proposed community-driven instruction engineering, where they collected instructions created by various NLP experts with different writing styles, diversifying the choices of instructions. However, human instruction engineering is time-consuming and expensive. Moreover, the human intuition on instruction designing might be subjective and sometimes is not optimal for the models.

### v8

@@ 5.1 Semantic Parser
At the early stage of machine learning, to help the systems understand natural language instructions, a great number of works employed semantic parsing to convert the instruction into the formal language (logical formula), which can be more easily executed by the systems (Goldwasser and Roth, 2011, inter alia). As exemplified in Figure 3 (a), a game instruction “Move any top card to an empty free cell” can be processed into an executable formula: “card(x) $\wedge$ freecell(y)”.
@@ 5.2 Flatten-and-Concatenation
In contrast to the semantic parser approach, which considers the instructions’ structure and the target problems, methods based on the neural networks take more brutal treatment: as illustrated in Figure 3 (b)—instructions, regardless of their length, structure, task types, etc., are flattened as a long token sequence and concatenated with the input X as a new input sequence for the models, which has been widely adopted by the prior research (Wang et al., 2022b; Wei et al., 2023, inter alia). However, this naive strategy constantly results in unsatisfactory performances when using vanilla models Weller et al. (2020), leading to its reliance on large-scale instruction fine-tuning, well-known as “instruction tuning”.
@@ 5.3 HyperNetwork
Unlike the conventional modeling strategy that encodes the input sequence into the dense representation (i.e., language-to-representation), hypernetwork follows a language-to-parameter paradigm: as shown in Figure 3 (c), this scheme converts textual instruction into a block of model parameters that can be further plugged into the underlying models Ha et al. (2017); Houlsby et al. (2019); Jin et al. (2020). As a result, hypernetwork-based instruction modeling can better leverage the structured input sequence by encoding the instruction and task input separately (i.e., instruction-to-parameter, input-to-representation), achieving stronger generalization compared with the flatten-and-concatenation approach Ye and Ren (2021); Deb et al. (2022). It can also significantly improve inference efficiency, as concluded by recent works Ivison et al. (2023a).
@@ 5.4 Reinforcement Learning from Human Feedback
As shown in Figure 3 (d), given the initial LM and its copy (the target tuned LM), for each input sequence (e.g., instruction and task input), these two different LMs will generate two outputs. After getting the generation texts of the initial and tuned LM, we can further calculate the textual difference penalty between them:
$y$ $=\theta(I,x)$
$y^{*}$ $=\theta^{*}(I,x)$
@@
Here, the $\theta$ and $\theta^{*}$ represent the parameters of initial and tuned LM, respectively. $I$ and $x$ denote the instruction and task input; While $y$ and $y^{*}$ are the outputs of the initial and tuned LM. $r_{\textrm{KL}}$ is the final reward (loss) of prediction shifting, and the KL means the calculation of Kullback–Leibler (KL) divergence. KL divergence is a widely adopted strategy for measuring the textual difference, which can be used as a part of the loss to penalize the tuned LM on shifting the output substantially away from the initial LM’s generation. This prediction shift penalty can prevent the tuned LM from fooling the reward function to get a high reward but loses the coherence of the generation text.
@@
The $r_{\theta}$ is the reward model, which is usually a regression model.
@@
Where the $\sigma$ is the activation function that scales the reward into $(0,1]$, $y^{+}$ is the output preferred by human, $y^{-}$ otherwise. By training on these pairwise preference comparison data, the reward model can directly learn to capture the human preference and make alignment reward estimation for the RLHF.
@@ 6.1 Dataset
The essence of instruction following is to tame the models by following various task instructions and responding with the corresponding desired outputs. Therefore, the instruction-tuning dataset (high-quality instruction-output pairs) is the critical part Wang et al. (2023c); Zhou et al. (2023a). The current instruction-tuning datasets can be divided into two categories according to different annotation categories: 1) human-annotated datasets (§ 6.1.1); 2) LLM-synthetic datasets (§ 6.1.2).
@@ 6.1.1 Human-annotated datasets
The conventional way to create instruction-tuning datasets is by employing human annotators, especially for those early-stage datasets, as shown in Table 6. For example, Public Pool of Prompts (P3) (Sanh et al., 2022) and Flan (Wei et al., 2022a) collected multi-task instruction-tuning datasets, where they employed human expertise to design various prompts for each task. Mishra et al. (2022b) proposed Natural Instructions, in which they collected more than 60 NLP tasks with the corresponding human-written instructions; Wang et al. (2022b) further extended this collection into a 1.6k cross-lingual tasks scale contributed by 88 NLP experts, namely Super-Natural Instructions.
@@ 6.1.2 LLM-synthetic datasets
Since LLMs have shown their superior annotation quality on various NLP tasks He et al. (2023); Pan et al. (2023), tons of recent works tried to employ LLMs (e.g., ChatGPT and GPT-4) instead of humans on instruction-tuning dataset curation. For instance, Self-Insturct Wang et al. (2023d) and Unnatural Instructions Honovich et al. (2023a) utilized human-annotated instructions as demonstrations to guide LLMs in devising novel tasks and increasing task diversity. WizardLM Xu et al. (2023a) employed an instruction evolution paradigm to increase instruction complexity. Dynosaur Yin et al. (2023) repurposed existing input-output pairs in NLP datasets to stimulate new instructions and reduce annotation costs. MUFFIN Lou et al. (2024) prompted the LLMs to gather different task instructions for the same input and obtained an impressive generalization capacity of the tuned smaller models. Besides single-turn instruction-output datasets, some works also collected multi-turn dialogue data from ShareGPT [^3] https://sharegpt.com/, where the instructions are created by humans (users of OpenAI API), and the responses are from LLMs Geng et al. (2023); Chiang et al. (2023).
Though these LLM-synthetic datasets contained considerable noise (e.g., incoherent instructions and hallucination outputs), the diverse task distribution and model-preferred output patterns still benefit the smaller models on instruction-following, achieving comparable even better generalization performance compared with human-annotated datasets Wang et al. (2023d, c).
In a word, the choice between human-annotated and LLM-synthetic datasets can also be regarded as a trade-off between data quality and diversity. Previous works have concluded that both factors affect the performance of the resulting models Chung et al. (2022); Longpre et al. (2023) — mixing human and machine data can lead to better results Wang et al. (2023d); Yin et al. (2023), while there is no concrete conclusion about which factor outweighs the other, which highly depends on the downstream tasks and application situations.
@@ 6.2.1 Different evaluation schemes
When testing the model’s instruction-following performance on an evaluation dataset, if this dataset has “ground-truth” outputs, then a conventional criterion is to use those automatic evaluation metrics, such as Exact-Match Rajpurkar et al. (2016) and ROUGE Lin (2004), that have been widely used for evaluating the generation models Mishra et al. (2022b); Wang et al. (2022b); Wei et al. (2022a); Sanh et al. (2022); Lou and Yin (2023); Yin et al. (2022). However, this naive evaluation strategy suffers from several drawbacks: 1) it has been widely committed that the automatic generation metrics are not perfect and have significant biases (e.g., BLUE score has text length bias); 2) all of these metrics are used for showing how well the model’s prediction aligns with pre-annotated answers, however, most real-world user tasks are highly open-ended, and there are probably no official ground-truth labels to calculate the metrics; 3) the essence of instruction following is to follow user’s instructions and provide desired responses that can well address user’s requirements, while automatic metrics more focus on some superficial textual patterns and lacking the reflection on how well the response satisfies the instructions.
@@
A more reliable evaluation method is to employ humans to decide whether a model’s response satisfies the instruction or not. For example, given a task instruction and a corresponding model output, the human evaluator should read the instruction and decide whether this model output is acceptable or not (reporting an acceptance ratio for the target model) Wang et al. (2023d, c); Lou et al. (2024); or ask humans to compare two models’ outputs and decide which one better satisfies the instruction (pairwise comparison between two models) Chiang et al. (2023); Geng et al. (2023); Taori et al. (2023). Since instructions are mostly complicated and contain a lot of explicit or implicit constraints, human evaluation is more flexible and accurate than automatic metrics in reflecting the instruction-following capacities of different models.
However, human evaluation is also much more expensive, slower than automatic evaluation, and unreproducible. Thus, most of the works only conduct human evaluation on a small subset of the whole evaluation benchmark. Meanwhile, it is mostly based on human evaluators’ personal preferences and can result in high variance between different evaluators.
@@
To address the aforementioned issues of human evaluation, recent works also tried to use LLMs (e.g., GPT-4) rather than humans to evaluate the models’ instruction following capacity, such as VicunaEval Chiang et al. (2023) and AlpacaEval Taori et al. (2023). Nevertheless, although LLMs are cheaper and faster, they were found to have serious preference bias on some superficial textual patterns or hallucinations, e.g., GPT-4 prefers the longer texts and those responses with diverse tokens Wang et al. (2023c). Meanwhile, only a final preference score is usually insufficient for a comprehensive evaluation.
In order to improve reliability, instead of letting LLMs simply provide a preference decision, other works tend to ask LLMs to generate comprehensive analyses besides the final decision, such as generating the error types, locations and explanations before concluding with the final scores Fernandes et al. (2023); Xu et al. (2023e). Some other works also predefined several explainable criterion questions for the various evaluation tasks (e.g., for an instruction “Please generate at least 25 sentences”, define a criterion “is the model’s generation at least 25 sentences?”), that can be further verified by humans or LLMs easily (i.e., doing binary classification on these predefined criteria) Liu et al. (2023e); Zhou et al. (2023b). Saha et al. (2023) also asked LLMs to first generate the criteria questions automatically according to the instructions and then evaluate the model’s response.
@@ 6.2.2 Two branches of evaluation
Most evaluation datasets in this branch are based on conventional multi-task learning, where the evaluation tasks are mostly traditional NLP tasks, such as natural language inference Wei et al. (2022a); Sanh et al. (2022). This branch aims to test LLMs’ instruction-following and problem-solving capacity, and the main criterion here is whether the models can correctly solve the given textual task. Therefore, most of the evaluation settings in this branch adopt conventional automatic metrics to reflect the task ground-truth label alignment. Representative benchmarks are MMLU Hendrycks et al. (2021), BBH Suzgun et al. (2023), SuperNI-Test Wang et al. (2022b), T0-Eval Sanh et al. (2022), InstructEval Chia et al. (2023), etc.
@@ 7 Factors that Influence Instruction Following Performance
{tblr}
[…]
Model-related Factors (§ 7.1)
@@
Instruction-related Factors (§ 7.2)
@@
Demonstration-related Factors(§ 7.3)
@@
Model-Instruction Alignment (§ 7.4)
@@
Data-wise Factors (§ 7.5)
@@ 7.1.1 Update model or not
As shown in Figure 1 (b), to drive LLMs to understand and follow task instructions more smoothly, a widely-adopted practice is fine-tuning LLMs on multi-task datasets, where each task input is equipped with a task instruction. This procedure is also well-known as “instruction tuning”. A lot of works demonstrated that instruction-tuned LLMs could better follow the instructions of unseen tasks compared with frozen LLMs Wei et al. (2022a); Sanh et al. (2022).
Besides the performance gains on unseen tasks, instruction tuning has many other benefits, such as learning faster on the downstream tasks Longpre et al. (2023); Gupta et al. (2023), being more robust to the tiny instruction perturbations (e.g., paraphrasing) Weller et al. (2020); Sanh et al. (2022); Gu et al. (2023), becoming more user-friendly Chung et al. (2022), and being better at following soft instructions Wei et al. (2022a), etc.
@@ 7.1.2 Model scale
Recent works demonstrated that the model scale significantly impacts the generalization performance of instruction following (Chung et al., 2022; Longpre et al., 2023; Wang et al., 2023c, inter alia). As shown in Figure 4, the generalization performances of each model consistently increase when scaling up the model size. More interestingly, when the model scale is large enough, even vanilla LLMs can significantly outperform smaller LLMs tuned on extensive tasks (see Flan-PaLM; vanilla 540B $>$ 8B + 1836 tasks), which probably implies that the benefits of scaling up the model size can outweigh dataset scaling.
However, the super-large model scale is usually unaffordable for most research groups, and it also leads to huge carbon emissions, making it unrealistic in most real-world scenarios Strubell et al. (2019); Schick and Schütze (2021c). Accordingly, recent works began to investigate a more efficient way to address the model scale problem, e.g., the parameter-efficient fine-tuning Hu et al. (2022a); Liu et al. (2022a); Lialin et al. (2023); Jang et al. (2023).
@@ 7.2.1 Instruction engineering
One straightforward solution is to manually rewrite the instruction, i.e., human instruction engineering. When humans perform instruction engineering, the criteria of rewriting are mostly human intuition. For example, Mishra et al. (2022a) conducted error cases analysis on GPT’s instruction-following outputs. Accordingly, they designed several empirical rules on instruction writing and “reframed” the instructions. All of these proposed rules are based on human intuition, e.g., itemizing instructions and task decomposition. In order to avoid the preference bias introduced by a small group of humans, Bach et al. (2022) proposed community-driven instruction engineering, where they collected instructions created by various NLP experts with different writing styles, diversifying the choices of instructions. However, human instruction engineering is time-consuming and expensive. Moreover, the human intuition on instruction designing might be subjective and sometimes sub-optimal for the models.